OpenStudy (anonymous):

I need some help with a question that was in my previous test, the teacher said that I was on the right track but not quite there. Find two consecutive odd integers such that five times their sum is 23 less than their product.

OpenStudy (faiqraees):

Let odd number first = x odd number second = x+2 5 times the sum + 23 = product of numbers 5(x+(x+2))+23 = x(x+2) Solve the equation for x

OpenStudy (anonymous):

thanks ill get down to it.